Solve each system by elimination. Indicate the POI.

A. -x - 2y = -1

B. x + 3y = -2
1
Expert's answer
2020-11-02T20:01:49-0500
"\begin{matrix}
 -x-2y=-1 \\
 x+3y=-2
\end{matrix}"

"\begin{matrix}
 x=-2y+1 \\
 -2y+1+3y=-2
\end{matrix}"

"\begin{matrix}
 x=-2y+1 \\
 y=-3
\end{matrix}"

"\begin{matrix}
 x=-2(-3)+1 \\
 y=-3
\end{matrix}"

"\begin{matrix}
 x=7 \\
 y=-3
\end{matrix}"

Point of intersection "(7, -3)."